76. Simplify.

Sqrt(6) Sqrt(14)
*
Sqrt(7)Sqrt(3)

114. Area of a triangle. Find the exact area of a triangle with
a base of Sqrt(30) meters and a height of Sqrt(6) meters.

Answer by edjones(8007)

sqrt(6*14)
=2sqrt(21)
.
sqrt(7*3)=sqrt(21)
.
2sqrt(21)*sqrt(21)
=2*21
=42
.
.
.5ab=A
.5*sqrt(30)*sqrt(6)
=.5sqrt(180)
=.5sqrt(36*5)
=.5*6sqrt(5)
=3sqrt(5)m^2
.
Ed